A mother is twenty years older than her daughter. Five years from now, she will be thrice as old as her daughter will then be. How old is the daughter now?
Answer by VFBundy(438)

Daughter's age now = d
Mother's age now = d + 20

(d + 20) + 5 = 3(d + 5)

d + 25 = 3d + 15

-2d = -10

d = 5

Daughter's age now = d = 5